# GMValuator: Similarity-based Data Valuation for Generative Models

### v7

@@ Abstract
Data valuation plays a crucial role in machine learning. Existing data valuation methods have primarily focused on discriminative models, neglecting generative models that have recently gained considerable attention. A very few existing attempts of data valuation method designed for deep generative models either concentrates on specific models or lacks robustness in their outcomes. Moreover, efficiency still reveals vulnerable shortcomings. To bridge the gaps, we formulate the data valuation problem in generative models from a similarity-matching perspective. Specifically, we introduce Generative Model Valuator (GMValuator), the first training-free and model-agnostic approach to provide data valuation for generation tasks. It empowers efficient data valuation through our innovatively similarity matching module, calibrates biased contribution by incorporating image quality assessment, and attributes credits to all training samples based on their contributions to the generated samples. Additionally, we introduce four evaluation criteria for assessing data valuation methods in generative models, aligning with principles of plausibility and truthfulness. GMValuator is extensively evaluated on various datasets and generative architectures to demonstrate its effectiveness.
I Introduction
@@
As the driving force behind modern artificial intelligence, particularly deep learning [1], a substantial volume of data is indispensable for effective machine learning. On one hand, informative data samples that are relevant to the task at hand play a critical role in the training process. On the other hand, due to data privacy concerns, personal data is safeguarded by various regulations, including the General Data Protection Regulation (GDPR) [2], and has become a valuable asset. Consequently, data valuation has garnered significant attention from both academic and industrial sectors in recent times.
@@
The intricate relationship between data and model parameters presents a significant challenge in contribution measurement of each training sample, thus making data valuation a difficult task. Most of the existing data valuation studies focus on supervised learning for discriminative models (e.g., classification and regression). These methods can be categorized as: (1) Metric-based methods: methods such as Shapley Value (SV) and Banzhaf Index(BI) [3, 4] provide the assessment of data value by calculating marginal contribution on performance metrics (e.g., accuracy or loss) through retraining the model [^1] Exception for KNN-Shap [5]. (2) Influence-based methods: this line of methods measures data value by evaluating influence on model parameters of data points [6, 7]. (3) Data-driven methods: These techniques avoid retraining by leveraging data characteristics (like data diversity, generalization bound estimation, class-wise distance), though generally necessitate data labels. [8, 9, 10].
@@
Data valuation in the context of generative models has NOT been well-investigated in the current literature. Moreover, there exist significant challenges in directly adapting the aforementioned data valuation methods for discriminative models listed to generative models: Firstly, the challenge of applying metric-based methods arises from lacking robust performance metrics in generative models, in contrast to the existance of commonly used metrics (e.g., accuracy or loss) in discriminative models. In addition, the expensive cost of retrain requirements is another obstacle to use metric-based methods. Secondly, influence-based methods may not perform well on non-convex objective function of generative models [11]. Besides, estimating influence function in deep generative model is expensive, as it requires computing (or approximation) inverse Hessian. Thirdly, for data-driven methods, they mainly focus on supervised learning, requires knowing data labels to quantify data values.
@@
To the best of our knowledge, limited studies have explored model-dependent data evaluation using influence functions for specific generative models. IF4GAN [12] consider multiple evaluation metrics, such as log-likelihood, inception score (IS), and Frechet inception distance (FID) [13] to identify the most responsible training samples for the overall performance of generative adversarial networks (GAN)-based models. However, the selection of appropriate metrics is critical, as the results are inconsistent across metrics. VAE-TracIn [14] finds the most significant contributors for generating a particular generated sample for Variational Autoencoders (VAE) using influence function. However, viable Hessian estimation in influence function calculations incur high computational costs and this method cannot be easily generalized to other generative models.
@@
Considering the challenges posed by the selection of performance metrics and computational efficiency, as well as the need for broader applicability across various generative models, we aim to propose a unified and efficient data valuation method. We expect the unified method to satisfy the following key properties: 1) Model-agnostic: the method should be versatile, capable of being employed across diverse generative model architectures and algorithms, regardless of specific design choices; 2) Computation Efficiency: the method does not require retraining the model and minimize computational overhead while maintaining a satisfactory level of accuracy and reliability in evaluating the value of data points; 3) Plausibility: the method should evaluate the value of data based on its alignment with human prior knowledge on the task, enhancing its credibility and reliability; 4) Truthfulness: the method should strive to provide an unbiased and accurate assessment of the value associated with individual data points. To meet the design purposes, we propose a similarity-based data valuation approach for the generative model, called GMValuator, which is model-agnostic and efficient.
@@
  • To the best of our knowledge, GMValuator is the first modal-agnostic and retraining-free data valuation method for generative models.
@@
  • We propose four evaluation methods to assess the truthfulness of data valuation and evaluate GMValuator on different datasets and deep generative models to verify its validity.
II Related Work
II-A Data Valuation
There are three lines of methods on data valuation: metric-based methods, influence-based methods and data-driven methods. In terms of metric-based methods, the commonly-used approach is to calculate its marginal contribution (MC) based on performance metrics (e.g., accuracy, loss). As the basic method depending on performance metrics for data valuation, LOO (Leave-One-Out) [15] is used to evaluate the value of training sample by observational change of model performance when leaving out that data point from the training dataset. To overcome inaccuracy and strict desirability of LOO, SV [3] and BI [4] originated from Cooperative Game Theory are widely used to measure the contribution of data [6, 16]. Considering the joining sequence of each training data point, SV needs to calculate the marginal performance of all possible subsets in which the time complexity is exponential. Despite the introduction of techniques such as Monte-Carlo and gradient-based methods, as well as others proposed in the literature, to approximate data significance value (SV) is computationally expensive and it typically require retraining [3, 5]. The computational cost and needs for unconventional performance metrics presents difficulties in adapting the methods to generative models. As for influence-based methods, they evaluate the influence of data points on model parameters by computing the inverse Hessian for data valuation [5, 17, 18]. Due to the high computational cost, some approximation methods have also been proposed [19]. In addition, the use of influence function for data valuation is not limited to discriminative models, but can also be applied to specific generative models such as GAN and VAE [12, 14]. When it comes to data-driven methods, most of them are training-free methods that focus on the data itself [8, 9, 10].
II-B Generative Model
Generative models are a type of unsupervised learning that can learn data distributions. Recently, there has been significant interest in combining generative models with neural networks to create Deep Generative Models, which are particularly useful for complex, high-dimensional data distributions. They can approximate the likelihood of each observation and generate new synthetic data by incorporating variations. Variational auto-encoders (VAEs) [20] optimize the log-likelihood of data by maximizing the evidence lower bound (ELBO), while generative adversarial networks (GANs) [21, 22] involves a generator and discriminator that compete with each other, resulting in strong image generation. Recently proposed diffusion models [23, 24] add Gaussian noise to training data and learn to recover the original data. These models use variational inference and have a fixed procedure with a high-dimensional latent space.
III Motivation and Problem Formulation
@@
In this section, we present the background and motivation behind our proposed method GMValuator, which formulates the data valuation for generative models as a similarity matching problem.
@@
III-A Motivation
Let $X=\left\{x_{1},...,x_{n}\right\}$ denote the training dataset for generative model training, in which $x\sim\mathcal{X}$ and $n$ is regarded as the size of data. Let $\mathcal{A}$ present the learning algorithm of generative models (e.g., Generative Adversarial Network (GAN), Variational Autoencoder (VAE), Diffusion Model). The primary goal of generative models is to generate $m$ synthetic data samples $\hat{X}=\left\{\hat{x}_{1},...,\hat{x}_{m}\right\}$ that are closely similar to the real data. Thus, the core of the generative model is an intractable probability distribution learning process to train a generator $\mathcal{A}$, that maps each sample $z$ sampled from the noise distribution $\mathcal{Z}$ to a sample $\hat{x}$ in estimated distribution $\mathcal{\hat{X}}$ while minimizing the discrepancy ($D$) between the estimated distribution $\mathcal{\hat{X}}$ and real data distribution $\mathcal{{X}}$:
$\mathcal{A}=\operatorname*{arg\,min}_{\mathcal{A}}D(\mathcal{A}_{z\sim\mathcal{Z}}(z)),\mathcal{{X}}),\quad\mathcal{A}:z\mapsto\hat{x}\sim\mathcal{\hat{X}}$
It’s important to understand the significance of each training data point in relation to the generated data. The objective function can be calculated using likelihood $p_{\mathcal{A}}(x|z)$ in (1), where $p_{\mathcal{A}}(x)$ qualifies the similarity between $\mathcal{A}(\mathbf{z})$ and $x$. The value of $p_{\mathcal{A}}$ should be high as $\hat{x}$ looks like the sampling data $x$, by maximizing the likelihood $p_{\mathcal{A}}(x|z)$.
[…]
Thus, generated data should exhibit a higher degree of similarity to the data points used to train the generator than the ones not used for training, despite originating from a similar distribution.
@@ III-A Motivation
Therefore, we put forth the plausible assumption that the similarity between training data and generated data characterizes the contribution of the training data to the generated data for any given model. We support the assumption by partitioning a class of CIFAR-10 (the class is plane here) into two non-overlapped subsets, denoted as $X_{v1}$ and $X_{v2}$. [^2] In practice, we perform T-SNE first to randomly split the non-overlapped samples from the embedding space. Next, we keep $X_{v1}$ as non-training data and use $X_{v2}$ as training data to train a BigGAN [25] to generate dataset $\hat{X}$. If our assumption holds, the generated data will be more similar to the training data $X_{v2}$. As presented in the T-SNE plot (Figure 1), the generated data demonstrate a more substantial overlap in the shared feature space with the training dataset $X_{v2}$ than the non-training ones $X_{v1}$. [^3] Quantitative statistical testing is provided in Appendix. This observation motivates us to address the data valuation for the generative model from a similarity-matching perspective.
It’s worth noting that concurrent to our research, LAVA [10] also proposes the utilization of a similarity metric, specifically the class-wise Wasserstein distance between the training and validation sets, to evaluate data value. However, their emphasis lies in classification tasks and theoretical analysis, whereas our work contributes practical algorithms tailored for generative models.
@@
III-B Problem Formulation and Challenges
The primary objective of our research is to tackle the issue of data valuation in generative models using black-box access. In essence, given a fixed set of data samples $\hat{X}$ with a size of $m$ generated from the trained deep generative model $\mathcal{A}^{*}$, our aim is to determine the value $\phi_{i}(x_{i},\hat{X},\mathcal{A}^{*})$ associated with each data point $x_{i}\in X$ for $i\in[n]$ in the deduplicated training dataset, which contributes to the generated dataset. We denote the value for training data $x_{i}$ as $\phi_{i}$ in the rest of the paper for simplicity.
Following the motivation stated in Sec. III-A, $\phi_{i}$ should be a functions of the distance between the data points. We denote the distance between training data $x_{i}$ and generated data $\hat{x}_{j}$ as $d_{ij}$, for $i\in[n]$ and $j\in[m]$. The contribution score of $x_{i}$ to $\hat{x}_{j}$ is denoted as $\mathcal{V}(x_{i},\hat{x}_{j})\propto d_{ij}^{-1}$, which is inversely proportional to distance since maximizing the log-likelihood of a generative model is equivalent to minimizing the dissimilarity of real and generated data distribution. To link the dissmilarity $-d_{ij}$ to likelihood, we choose $exp(-d_{ij})$ to likelihood, and define $\mathcal{V}(x_{i},\hat{x}_{j})$ as follows:
Therefore, an intuitive definition for the value of data $i$ can be written as
$\phi_{i}=\sum_{j=1}^{m}\mathcal{V}(x_{i},\hat{x}_{j}),x_{i}\in X,\hat{x}_{j}\in\hat{X}.$ (3)
However, there are several open questions and challenges in performing the above two steps for calculating Eq. (3):
Challenge 1: Efficiency. In step 1, considering $n$ training samples and $m$ generated samples, where $\mathcal{O}(C)$ represents the complexity of the selected pair-wise distance calculation, the total complexity of this step amounts to $\mathcal{O}(mnC)$. In practical scenarios with large training datasets (e.g., $n>10K$), the computation cost for becomes prohibitively expensive. Additionally, fitting such a large collection of high-dimensional data for distance calculation can pose significance challenges in system memory.
Challenge 3: Non-zero scores. In practical scenarios, the distance between training data and the least similar generated data is not infinite, which may result in false non-zero contribution scores. With a large dataset size, the accumulation of these noisy scores yield biased data valuation.
In light of this, we present a novel and efficient data valuation approach suitable for agnostic generative models, termed as GMValuator as elaborated in Section IV.
IV The Proposed Data Valuation Methods

### v8

@@ Abstract
Data valuation plays a crucial role in machine learning. Existing data valuation methods, mainly focused on discriminative models, overlook generative models that have gained attention recently. In generative models, data valuation measures the impact of training data on generated datasets. Very few existing attempts at data valuation methods designed for deep generative models either concentrate on specific models or lack robustness in their outcomes. Moreover, efficiency still reveals vulnerable shortcomings. We formulate the data valuation problem in generative models from a similarity matching perspective to bridge the gaps. Specifically, we introduce Generative Model Valuator (GMValuator), the first training-free and model-agnostic approach to providing data valuation for image generation tasks. It empowers efficient data valuation through our innovative similarity matching module, calibrates biased contributions by incorporating image quality assessment, and attributes credits to all training samples based on their contributions to the generated samples. Additionally, we introduce four evaluation criteria for assessing data valuation methods in generative models. GMValuator is extensively evaluated on benchmark and high-resolution datasets and various mainstream generative architectures to demonstrate its effectiveness. Our code is available at: https://github.com/ubc-tea/GMValuator.
1 Introduction
@@
As the driving force behind modern AI, particularly deep learning Pei (2020), a substantial volume of data is indispensable for effective machine learning. On one hand, informative data samples relevant to the task at hand play a critical role in the training process. On the other hand, due to data privacy concerns, personal data is safeguarded by various regulations, including the General Data Protection Regulation (GDPR) Regulation (2018), and has become a valuable asset. Consequently, data valuation has garnered significant attention from academic and industrial sectors recently.
@@
The intricate relationship between data and model parameters presents a significant challenge in the contribution measurement of each training sample, thus making data valuation a difficult task. Most existing data valuation studies focus on supervised learning for discriminative models (e.g., classification and regression). These methods can be categorized as (1) Metric-based methods: Methods such as Shapley Value (SV) and Banzhaf Index (BI) Ghorbani & Zou (2019); Wang & Jia (2023) provide the assessment of data value by calculating marginal contribution on performance metrics (e.g., accuracy or loss) through retraining the model [^1] Exception for KNN-Shap Jia et al. (2019a). (2) Influence-based methods: This line of methods measures data value by evaluating influence on model parameters of data points Jia et al. (2019b); Nohyun et al. (2022). (3) Data-driven methods: These techniques avoid retraining by leveraging data characteristics (like data diversity, generalization bound estimation, class-wise distance), though generally necessitate data labels. Xu et al. (2021); Wu et al. (2022); Just et al. (2023).
@@
Urgent needs for generative models. Data valuation in the context of generative models has NOT been well-investigated in the current literature. Moreover, there exist significant challenges in directly adapting the aforementioned data valuation methods for discriminative models listed to generative models: Firstly, the challenge of applying metric-based methods arises from lacking robust performance metrics in generative models, in contrast to the existence of commonly used metrics (e.g., accuracy or loss) in discriminative models Betzalel et al. (2022). This could potentially result in inconsistent outcomes when employing various performance metrics Terashita et al. (2021). In addition, the expensive cost of retraining requirements is another obstacle to using metric-based methods. Secondly, influence-based methods may not perform well on non-convex objective function of generative models Bae et al. (2022). Besides, estimating the influence function in a deep generative model is expensive, as it requires computing (or approximation) inverse Hessian. Thirdly, for data-driven methods, they mainly focus on supervised learning, which requires knowing data labels to quantify data values.
@@
To the best of our knowledge, limited studies have explored model-dependent data evaluation using influence functions for specific generative models. IF4GAN Terashita et al. (2021) consider multiple evaluation metrics, such as log-likelihood, Inception Score (IS), and Frechet Inception Distance (FID) Heusel et al. (2017) to identify the most responsible training samples for the overall performance of Generative Adversarial Networks (GAN)-based models. However, selecting appropriate metrics is critical, as the results are inconsistent across metrics. VAE-TracIn Kong & Chaudhuri (2021) finds the most significant contributors for generating a particular generated sample for Variational Autoencoders (VAE) using the influence function. However, viable Hessian estimation in influence function calculations incur high computational costs and this method cannot be easily generalized to other generative models.
@@
Our motivation and goal. Considering the challenges posed by the selection of performance metrics and computational efficiency and the need for broader applicability across various generative models, we aim to propose a unified and efficient data valuation method. We expect the unified method to satisfy the following key properties: 1) Model-agnostic: the method should be versatile, capable of being employed across diverse generative model architectures and algorithms, regardless of specific design choices; 2) Computation Efficiency: the method does not require retraining the model and minimize computational overhead while maintaining a satisfactory level of accuracy and reliability in evaluating the value of data points; 3) Plausibility: the method should evaluate the value of data based on its alignment with human prior knowledge on the task, enhancing its credibility and reliability; 4) Truthfulness: the method should strive to provide an unbiased and accurate assessment of the value associated with individual data points. In this work, we evaluate the contribution of (good) training data given a fixed set of generated data from a certain well-trained generative model. We refer to the value of training data as the contribution to the given generated data, taking into account the quality of the generated data. To meet the design purposes, we propose a similarity-based data valuation approach   [^2] We detail why the similarity-based approach is formed in Sec 2 and explain how it is implemented in Sec 3. for the generative model, called Generative Model Valuator (GMValuator), which is model-agnostic and efficient.
@@
• To the best of our knowledge, GMValuator is the first modal-agnostic and retraining-free data valuation method for image generative models.
@@
• We propose four evaluation methods to assess the truthfulness of data valuation and evaluate GMValuator on different datasets (including benchmark datasets and high-resolution large-scale datasets) and various deep generative models to verify GMValuator’s validity.
Related Work. Different from discriminative models for regression or classification tasks, generative models in various forms (i.e., Variational auto-encoders (VAEs) Rezende et al. (2014), Variational auto-encoders (VAEs) Rezende et al. (2014), Diffusion Model Ho et al. (2020); Rombach et al. (2022)) aim to learn data distribution for generation tasks. Consequently, the approach to data valuation in our work, which focuses on generative models, is distinct and orthogonal to that for discriminative models. The primary limitation of both metric-based methods and influence-based methods is the expensive computation cost, a drawback that is further magnified in generative models. Conversely, while data-driven methods are training-free, they predominantly concentrate on supervised learning and data itself Xu et al. (2021); Wu et al. (2022); Just et al. (2023). Apart from the expensive computational cost, the limited existing influence-based methods for generative models are model-specific, which can not adapt to the diverse and evolving trends in generative model development Terashita et al. (2021); Kong & Chaudhuri (2021). A more detailed related work is discussed in Sec. C of the appendix.
2 Motivation and Problem Formulation
@@
In this section, we present the background and motivation behind our proposed method GMValuator, which formulates the data valuation for generative models as a similarity-matching problem.
2.1 Motivations
Let us denote $X=\left\{x_{1},...,x_{n}\right|x\sim\mathcal{X}\}$ as the training dataset without duplicated data for a generation task, and $n$ is regarded as the size of the dataset. Denote the generated dataset as $\hat{X}=\{\hat{x}_{1},...,\hat{x}_{m}\}$ by a well trained generative model $G^{*}$ (e.g., Generative Adversarial Network (GAN), Variational Auto-encoder (VAE), Diffusion Model) on $X$.
[…]
$G^{*}=\operatorname*{arg\,min}_{G}d(G_{z\sim\mathcal{Z}}(z)),\mathcal{{X}}).$ (1)
Therefore, the data generated by an optimal generator, which closely approximates the training distribution, can be considered as drawing from a subdistribution of $\mathcal{X}$.
2.1.1 Theoretical Justification
Let $T\subset X$ denote the $K$ contributors found by a data valuator (e.g., GMValuator) for $\hat{X}$, and $S$ is a subset of the training dataset $X$. We focus on the data type with describable attributes (e.g., image data with semantic attributes). We first provide definitions of data and contributors.
Definition 2.1
(Data with Describable Attributes.) We assume an image can be characterized by $V$ attributes, and there is a labeling function $f$ mapping $S$, $T$ and $\hat{X}$ to the same attribute space $\mathcal{A}=\{0,1\}^{V}$.
Definition 2.2
(Contributors in Generative Models.) Let $S^{*}$ denotes the real $K$contributors, defined as follows:
$S^{*}=\operatorname*{arg\,min}_{S\subset X}d(\mathcal{X}_{(\hat{X}|{A})},\mathcal{X}_{(G(S)|{A})}),$ (2)
where $|S^{*}|=K$ and $K$ is a reasonable number for generative model training. $\mathcal{X}_{(\hat{X}|A)}$ is the data distribution of generated data $\hat{X}$ on attributes $A\in\mathcal{A}$, and $\mathcal{X}_{(G(S)|A)}$ is the distribution of data with attribute $A$ that are generated by the optimal generator trained by contributors $S$. According to the objective of generative models and Eq. equation 2, we have $\mathcal{X}_{(\hat{X}|A)}\sim\mathcal{X}_{(G(S^{*})|{A})}$. We follow Just et al. (2023) on assumptions and lemmas that will be used to obtain the theorem.
Assumption 2.3
Assume that the function $f$ is $\epsilon-$Lipschitz and the loss function :$\mathcal{L}:\{0,1\}^{V}\times[0,1]^{V}\rightarrow R^{+}$ is k-Lipschitz in both inputs and attributes. We have labeling functions that are all bounded by $V$ as $\|f\|\leq V$.
Then, we introduce the error bound between $T$ and $S^{*}$.
Theorem 2.4
(Bounded Attributes Classification Error on $S^{*}$ to $T$.) Let $f_{S^{*}}^{{}^{\prime}}:\mu\rightarrow\mathcal{A}=\{0,1\}^{V}$ be the model trained on the optimal contributor dataset $S^{*}$. Following Assumption 2.3, if the contributors are corresponding to the given generated data $\hat{X}$, we have: (Jiaxi: $S^{*}$ in the appendix)
$\mathbb{E}_{x\sim\mu_{T}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]-\mathbb{E}_{x\sim\mu_{S^{*}}}\left[\mathcal{L}\left(f(x),f_{S^{*}}^{{}^{\prime}}(x)\right)\right]$ (3)
$\leq$ $k\epsilon\cdot\!\left[d_{W}(\mathcal{X}_{(T|f)},\mathcal{X}_{(\hat{X}|f)})\!+\!d_{W}(\mathcal{X}_{(S^{*}|f)},\mathcal{X}_{(\hat{X}|{f})})\right]$
As shown in the Theorem 2.4, given the optimal contributor $S^{*}$, $d_{W}\left(\mathcal{X}_{(S^{*}|f)},\mathcal{X}_{(\hat{X}|{f})}\right)$ is a deterministic term that approaching zero. Then, approximate $S^{*}$ can be achieved by reducing the distance term $d_{W}\left(\mathcal{X}_{(T|f)},\mathcal{X}_{(\hat{X}|f)}\right)$. Please see Sec. M in the appendix for the proof.
2.1.2 Emperical Validation
@@
According to theoretical justification, the generated data should show more similarity to the data samples used for training. In this context, we examine the value of training data being more valuable compared to data that is not used for training, despite originating from a similar distribution. We support this by partitioning a class of CIFAR-10 (the class is plane here) into two non-overlapped subsets, denoted as $X_{v1}$ and $X_{v2}$. [^3] In practice, we perform T-SNE first to randomly split the non-overlapped samples from the embedding space. Next, we keep $X_{v1}$ as non-training data and use $X_{v2}$ as training data to train a BigGAN Brock et al. (2018) and generate dataset $\hat{X}$. If our assumption holds, the generated data will be more similar to the training data $X_{v2}$. As presented in the T-SNE Van der Maaten & Hinton (2008) plot (Figure 1), the generated data demonstrate a more substantial overlap in the shared feature space with the training dataset $X_{v2}$ than the non-training ones $X_{v1}$. [^4] Quantitative statistical testing is provided in Appendix. This observation motivates us to address the data valuation for the generative model from a similarity-matching perspective.
2.2 Problem Formulation and Challenges
The primary objective of our research is to tackle the issue of data valuation in generative models using black-box access. In essence, given a fixed set of data samples $\hat{X}$ with a size of $m$ generated from the well-trained deep generative model $G^{*}$, our aim is to determine the value $\phi_{i}(x_{i},\hat{X},G^{*})$ associated with each data point $x_{i}\in X$ for $i\in[n]$ in the deduplicated training dataset, which contributes to the generated dataset. We denote the value for training data $x_{i}$ as $\phi_{i}$ in the rest of the paper for simplicity.
Following the motivation stated in Sec. 2.1, $\phi_{i}$ should be a function of the distance between the data points. We denote the distance between training data $x_{i}$ and generated data $\hat{x}_{j}$ as $d_{ij}$, for $i\in[n]$ and $j\in[m]$.
Definition 2.5
(Primary Contribution Score.) The contribution score of $x_{i}$ to $\hat{x}_{j}$ is denoted as $\mathcal{V}(x_{i},\hat{x}_{j})\propto d_{ij}^{-1}$, which is inversely proportional to distance since maximizing the log-likelihood of a generative model is equivalent to minimizing the dissimilarity of real and generated data distribution. To link the dissimilarity $-d_{ij}$ to likelihood, we choose $exp(-d_{ij})$ to likelihood
Therefore, an intuitive definition for the value of data sample $x_{i}$ can be written as below.
Definition 2.6
(Data Value.) The contribution of each training data point $x_{i}\in S^{*}$ for the generation of dataset $\hat{X}$ equals to the sum of its contributions to each generated data point $\hat{x}_{i}$ in $\hat{X}$.
$\phi_{i}=\sum_{j=1}^{m}\mathcal{V}(x_{i},\hat{x}_{j}),x_{i}\in X,\hat{x}_{j}\in S^{*}.$ (5)
However, there are several open questions and challenges in performing the above two steps for calculating Eq. equation 5:
Challenge 1: Efficiency. In step 1, considering $n$ training samples and $m$ generated samples, where $\mathcal{O}(C)$ represents the complexity of the selected pair-wise distance calculation, the total complexity of this step amounts to $\mathcal{O}(mnC)$. In practical scenarios with large training datasets (e.g., $n>10K$), the computation cost becomes prohibitively expensive. Additionally, fitting such a large collection of high-dimensional data for distance calculation can pose significant challenges in system memory.
Challenge 3: Non-zero scores. In practical scenarios, the distance between training data and the least similar generated data is not infinite, which may result in false non-zero contribution scores. With a large dataset size, the accumulation of these noisy scores yields biased data valuation.
In light of this, we present a novel and efficient data valuation approach suitable for agnostic generative models, termed as GMValuator as elaborated in Section 3.
@@ 2.2 Problem Formulation and Challenges
3 The Proposed Data Valuation Methods